# The leading constant in Malle’s conjecture over function fields

### v1

@@ 1. Introduction
Let $G$ be a finite group. Based on numerical data Malle [Malle2002Galois, Malle2004Galois] made a prediction on the asymptotic distribution of $G$-extensions $\mathbb{Q}$ of bounded discriminant. These predictions have been generalized from $\mathbb{Q}$ to general global fields and to more general ways of enumerating $G$-extensions than the discriminant.
@@ 1.4. Notations and conventions
$\operatorname{Br}_{\mathcal{C},\ell}BG_{\mathbb{F}_{q}}$ Brauer elements whose residues are $\ell$.
@@ 2.1. Root stacks
Let $G$ be a finite étale group scheme over a base scheme $G$. The goal of this section will be to recall some of the theory relevant to the study of $G$-covers. If $G$ is constant group scheme then the notion of the cover of a smooth scheme is well-known. But in this paper $G$ will in general be non-constant, which is slightly subtler.
@@
Let $X$ be a smooth and finitely presented scheme over $B$. A left $G$-cover of $X$ is a finite map of $B$-stacks $f:Y\to X$, where $Y$ is equipped with a left $G$-action such that $f$ is a $G$-torsor when restricted to a $B$-dense open substack $U\subset X$. We call the complement of the maximal such $U$ the ramification divisor of $f$.
@@
If $C$ is proper and $B$ Noetherian then is [Cadman2007stacks, Thm. 4.1.]. The proof is local and never uses properness, the general case is reduced to the Noetherian case in the standard way using that $f$ is finitely presented. ∎
@@
Let $C$ be a curve over $B$ and $f:X\to C$ a $G$-cover. The ramification divisor $D$ of $f$ is a divisor and finite étale over $B$. Moreover, for any finite étale divisor $D\subset C$ the following three categories are equivalent, moreover the equivalence is functorial in $B$ and $G$.
@@
The compositions of these three functors are equivalent to the identity. Indeed all stacks are separated so their morphisms are determined by the open dense subscheme $C\setminus D$ where the functors are the identity. ∎
@@ 2.3. Ramification type
Moreover, the composition $(D_{i})_{S}\to\mathcal{C}_{G}^{*}$ descends to a map $D_{i}\to\mathcal{C}_{G}^{*}$ as it is a map of schemes and the chosen isomorphism $(\sqrt[w_{i}]{D_{i}})_{S}\cong(B\mu_{\underline{w}})_{S}$ is unique. We call the union of maps $D=\sqcup D_{i}\to\mathcal{C}_{G}^{*}$ map the ramification type and denote it by $\rho_{G,D}(\varphi)$. More generally, for any subdivisor $D^{\prime}\subset D$ we write $\rho_{G,D^{\prime}}(\varphi)=\rho_{G,D}|_{D^{\prime}}:D^{\prime}\to\mathcal{C}_{G}*$. If $B$ is the spectrum of a field $l$ and $P\in D(k)$ then $\rho_{G,P}(\psi)\in\mathcal{C}_{G}^{*}(R)$ is exactly the ramification type defined in [Loughran2025mallesconjecturebrauergroups, §7.1].
@@ 2.5. Twisted Hurwitz schemes
See (2.1) for why this chart makes sense.
Recall the notion of a $G$-cover from §2.2
@@ 2.6. Hurwitz stacks
The scheme $\mathrm{Hur}^{G,\varphi}_{B,n}\to BG$ is the fiber of the map $\mathcal{H}\mathrm{ur}^{G}_{B,n}\to BG$ at $\varphi\in BG(B)$.
@@
If $\varphi=e_{BG}$, this corresponds to the fact that the Hurwitz stack is $\mathrm{Hur}^{G}_{B,n}/G]$, c.f. [ellenberg2025homologicalstabilitygeneralizedhurwitz, Rem. 2.4.6]. This is the way also the way the Hurwitz stack is defined in [ellenberg2025homologicalstabilitygeneralizedhurwitz, Landesman2025Homological].
@@
For any conjugacy-invariant subscheme $C\subset G(-1)\setminus\{1\}$ we define $\mathrm{Hur}^{G,C,(\varphi,\gamma)}_{B,\underline{n}}:=\mathrm{Hur}^{G,(\varphi,\gamma)}_{B}\cap\mathrm{Hur}^{G,C,\varphi}_{B,\underline{n}}\subset\mathrm{Hur}^{G,\varphi}_{B}$ and $\mathrm{CHur}^{G,C,(\varphi,\gamma)}_{B,\underline{n}}:=\mathrm{Hur}^{G,(\varphi,\gamma)}_{B}\cap\mathrm{CHur}^{G,C,\varphi}_{B,\underline{n}}\subset\mathrm{Hur}^{G,\varphi}_{B}$.
Note that $\mathrm{Hur}^{G,(\varphi,\gamma)}_{B,n}$ is an open and closed substack by Lemma˜2.19 and because the map $G(-1)/G]\to BG$ is finite étale.
@@ 2.7. Functoriality
We have to show for every $G$-cover $X\to\mathbb{P}^{1}_{T}$ we can find a unique up to unique isomorphism $(D,i)$ as in the definition. This pair is determined uniquely up to unique isomorphism by its image, so this follows from Proposition˜2.4. ∎
@@
If $f(C_{G})\subset C_{H}$ then The natural map $\mathrm{Hur}^{G,\varphi}\to\mathrm{Hur}^{H,f_{*}(\varphi)}$ maps $\mathrm{Hur}^{G,C_{G},\varphi}$ into $\mathrm{Hur}^{H,C_{G},f_{*}(\varphi)}$.
@@ 3.2. Stable components in all directions
By working étale locally on $B$ we may assume that $G$ is constant, $\varphi$ is trivial and $\mu_{|G|}$ is a constant finite étale scheme. This case reduces by functoriality to the universal case $B=\operatorname{Spec}\mathbb{Z}[\mu_{|G|}\frac{1}{|G|}]$. In this case it is a map of ind-finite étale schemes so it suffices to prove it over one geometric point in each connected component, so we may assume that $B=\operatorname{Spec}\mathbb{C}$.
@@
There exists $N>0$ such that if $n_{c}\geq N$ for all $c\in\mathcal{C}_{\mathrm{st}}$ and $n_{c}>0$ for $c\not\in\mathcal{C}_{\mathrm{st}}$ then the map $\pi_{0}(\mathrm{CHur}^{G,(\varphi,\gamma)}_{B,\underline{n}}/B)\to U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}$ is an isomorphism.
@@ 3.3. Brauer group computations
We need to recall some notation from [Loughran2025mallesconjecturebrauergroups] to define this subgroup of the Brauer group. We have provided an alternative description of this subgroup in terms of certain central extensions in §3.4 for the reader unfamiliar with [Loughran2025mallesconjecturebrauergroups]. Such a reader may take this alternative description as a definition.
@@
3.4. Description of the Brauer group
We will use the non-standard notation $\mathbb{Q}/\mathbb{Z}(1):=\lim_{p\not\mid n}\mu_{n}$ to keep the notation light.
@@ 3.4. Description of the Brauer group
Consider a central extension of group schemes $1\to\mathbb{Q}/\mathbb{Z}(1)\to E\to G\to 1$ of group schemes over $\mathbb{F}_{q}$. This defines an element $E]\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ by [Loughran2025mallesconjecturebrauergroups, Def. 6.5]. This defines an isomorphism between $BG_{\bar{\mathbb{F}}_{q}}$ and the group of $\mathbb{Q}/\mathbb{Z}(1)$ central extensions up to isomorphism under the Baer sum by taking the limit in $n$ of [Loughran2025mallesconjecturebrauergroups, Lem. 6.6,6.7] as $\operatorname{Br}\mathbb{F}_{q}=0$.
@@
Both statements follows immediately from [Loughran2025mallesconjecturebrauergroups, Lem. 6.9] by taking the colimit as $n\to\infty$ and writing down the torsors in loc. cit. in terms of cocycles. ∎
3.5. A computation
To prove this theorem we will essentially need to do some computations in the derived category $\mathcal{D}(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$. But everything is simple enough that we can give cocycle-theoretic descriptions of all involved objects and will not actually need any facts about derived categories. We leave the argument that the cocycle-theoretic definitions agree with the analogous definitions in the derived category to the interested reader, it follows from a standard bar complex argument.
@@ 3.5. A computation
The expression $\psi(\mathrm{Frob}_{q})(\underline{n}+\mathfrak{R}(\tilde{\gamma}))-\alpha_{*}(\eta)(\mathrm{Frob}_{q}^{-1}))$ is equal to
@@ 4.1. Point-counting on Hurwitz spaces
    Consider a chain of subsets $\mathcal{C}_{\mathrm{st}}\subseteq\mathcal{C}\subseteq\mathcal{C}_{G}^{*}$ such that the conjugacy classes in $\mathcal{C}_{\mathrm{st}}$, resp. $\mathcal{C}$, generate $G/Z(G)$, resp. $G$.
@@
    $\mathrm{CHur}^{G,(\varphi,\gamma)}_{\underline{n}}(\mathbb{F}_{q})=\#U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}(\mathbb{F}_{q})\#\mathrm{Conf}_{\mathcal{C},\underline{n}}(\mathbb{F}_{q})+O\left(q^{|\underline{n}|-\delta\min_{c\in\mathcal{C}_{\mathrm{st}}(\bar{\mathbb{F}}_{q})}(n_{c})}\right).$ (4.2)
@@
Let us now consider (2). By part (1) and Lemma 3.23 we may assume that the Frobenius fixed components of $\mathrm{CHur}^{G,(\varphi,\gamma)}_{\underline{n}}$ are in bijection with $U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}(\mathbb{F}_{q})$. It thus suffices to show that if $Z\subset\mathrm{CHur}^{G,\varphi}_{\underline{n}}$ is a Frobenius fixed component of a Hurwitz scheme then
@@
Let $\mathcal{C}_{\mathrm{st}}\subset\mathcal{C}\subset\mathcal{C}_{G}^{*},\underline{n}$ be as in Lemma˜4.1(2). Then $\mathrm{CHur}_{\underline{n}}^{G,(\varphi,\gamma)}(\mathbb{F}_{q})$ is equal to
$\frac{\#\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})}{|G^{\mathrm{ab}}(-1)(\mathbb{F}_{q})|}\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}}e(\partial_{\gamma}(\beta)(\varphi))\prod_{c\in\pi_{0}(\mathcal{C})}e(n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta))+O(q^{|\underline{n}|-\delta\min_{c\in\mathcal{C}_{\mathrm{st}}(\bar{\mathbb{F}}_{q})}(n_{c})}).$
@@
$\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}}e\left(\partial_{\gamma}(\beta)(\varphi)+\sum_{c\in\pi_{0}(\mathcal{C})}n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta)))\right)\frac{\#\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})|\operatorname{H}_{2}(G,\mathcal{C})(\mathbb{F}_{q})|}{|\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}|}.$
@@ 4.2.2. Brauer-Manin pairing
If $P=\infty$ then we have the following explicit formula , where we recall from §3.3 that $\partial_{\gamma}(\beta)(\varphi):=\partial_{c_{\gamma}}(\beta)((\varphi,\gamma))$.

### v2

@@ 1. Introduction
Let $G$ be a finite group. Based on numerical data Malle [Malle2002Galois, Malle2004Galois] made a prediction on the asymptotic distribution of $G$-extensions of $\mathbb{Q}$ of bounded discriminant. These predictions have been generalized from $\mathbb{Q}$ to general global fields and to more general ways of enumerating $G$-extensions than the discriminant.
@@ 1.4. Notations and conventions
$\operatorname{Br}_{\mathcal{C},\ell}BG_{\mathbb{F}_{q}}$ Brauer elements whose residues are $\ell$. 3.26
@@ 2.1. Root stacks
Let $G$ be a finite étale group scheme over a base scheme $B$. The goal of this section will be to recall some of the theory relevant to the study of $G$-covers. If $G$ is constant group scheme then the notion of the cover of a smooth scheme is well-known. But in this paper $G$ will in general be non-constant, which is slightly subtler.
@@
Let $X$ be a smooth and finitely presented scheme over $B$. A left $G$-cover of $X$ is a finite map of $B$-schemes $f:Y\to X$, where $Y$ is equipped with a left $G$-action such that $f$ is a $G$-torsor when restricted to a $B$-dense open substack $U\subset X$. We call the complement of the maximal such $U$ the ramification divisor of $f$.
@@
If $C$ is proper and $B$ Noetherian then this is [Cadman2007stacks, Thm. 4.1.]. The proof is local and never uses properness, the general case is reduced to the Noetherian case in the standard way using that $f$ is finitely presented. ∎
@@
Let $C$ be a curve over $B$ and $f:X\to C$ a $G$-cover. The ramification divisor $D$ of $f$ is a divisor and finite étale over $B$. Moreover, for any finite étale divisor $D\subset C$ the following three categories are equivalent, moreover the equivalences are functorial in $B$ and $G$.
@@
Any composition of these three functors is equivalent to the identity. Indeed all stacks are separated so their morphisms are determined by the open dense subscheme $C\setminus D$ where the functors are the identity. ∎
@@ 2.3. Ramification type
Moreover, the composition $(D_{i})_{S}\to\mathcal{C}_{G}^{*}$ descends to a map $D_{i}\to\mathcal{C}_{G}^{*}$ as it is a map of schemes and the chosen isomorphism $(\sqrt[w_{i}]{D_{i}})_{S}\cong(B\mu_{\underline{w}})_{S}$ is unique. We call the union of maps $D=\bigsqcup_{i}D_{i}\to\mathcal{C}_{G}^{*}$ map the ramification type and denote it by $\rho_{G,D}(\varphi)$. More generally, for any subdivisor $D^{\prime}\subset D$ we write $\rho_{G,D^{\prime}}(\varphi)=\rho_{G,D}|_{D^{\prime}}:D^{\prime}\to\mathcal{C}_{G}^{*}$. If $B$ is the spectrum of a field $k$ and $P\in D(k)$ then $\rho_{G,P}(\psi)\in\mathcal{C}_{G}^{*}(R)$ is exactly the ramification type defined in [Loughran2025mallesconjecturebrauergroups, §7.1].
@@ 2.5. Twisted Hurwitz schemes
See (2.1) for this chart.
@@ 2.6. Hurwitz stacks
The scheme $\mathrm{Hur}^{G,\varphi}_{B,n}\to B$ is the fiber of the map $\mathcal{H}\mathrm{ur}^{G}_{B,n}\to BG$ at $\varphi\in BG(B)$.
@@
If $\varphi=e_{BG}$, this corresponds to the fact that the Hurwitz stack is $\mathrm{Hur}^{G}_{B,n}/G]$, c.f. [ellenberg2025homologicalstabilitygeneralizedhurwitz, Rem. 2.4.6]. This is also the way the Hurwitz stack is defined in [ellenberg2025homologicalstabilitygeneralizedhurwitz, Landesman2025Homological].
@@
For any conjugacy-invariant subscheme $C\subset G(-1)\setminus\{1\}$ we define $\mathrm{Hur}^{G,C,(\varphi,\gamma)}_{B,\underline{n}}:=\mathrm{Hur}^{G,(\varphi,\gamma)}_{B}\cap\mathrm{Hur}^{G,C,\varphi}_{B,\underline{n}}\subset\mathrm{Hur}^{G,\varphi}_{B}$ and $\mathrm{CHur}^{G,C,(\varphi,\gamma)}_{B,\underline{n}}:=\mathrm{Hur}^{G,(\varphi,\gamma)}_{B}\cap\mathrm{CHur}^{G,C,\varphi}_{B,\underline{n}}\subset\mathrm{Hur}^{G,\varphi}_{B}$. The definition implies that all squares in the following diagram are cartesian
[…]
Note that $\mathrm{Hur}^{G,C,(\varphi,\gamma)}_{B,n}\subset\mathrm{Hur}^{G,C,\varphi}_{B,\underline{n}}$ is an open and closed substack by Lemma˜2.19 and because the map $B\xrightarrow{\gamma}G_{\varphi}(-1)$ is finite étale.
@@ 2.7. Functoriality
This will follow from Proposition˜2.4. Given that proposition it remains to show that for every $G$-cover $X\to\mathbb{P}^{1}_{T}$ that we can find a unique up to unique isomorphism pair $(D,i)$ as in the definition. This is true because such a pair is determined uniquely up to unique isomorphism by its image which we get from Proposition˜2.4. ∎
@@
The natural map $\mathrm{Hur}^{G,\varphi}\to\mathrm{Hur}^{H,f_{*}(\varphi)}$ maps $\mathrm{Hur}^{G,C_{G},\varphi}$ into $\mathrm{Hur}^{H,C_{G},f_{*}(\varphi)}$.
@@ 3.2. Stable components in all directions
By working étale locally on $B$ we may assume that $G$ is constant, $\varphi$ is trivial and $\mu_{|G|}$ is a constant finite étale scheme. This case reduces by functoriality to the universal case $B=\operatorname{Spec}\mathbb{Z}[\mu_{|G|},\frac{1}{|G|}]$. In this case it is a map of ind-finite étale schemes so it suffices to prove it over one geometric point in each connected component, so we may assume that $B=\operatorname{Spec}\mathbb{C}$.
@@
There exists $N>0$ such that if $n_{c}\geq N$ for all $c\in\mathcal{C}_{\mathrm{st}}$ and $n_{c}>0$ for $c\not\in\mathcal{C}_{\mathrm{st}}$ then the map $\pi_{0}(\mathrm{CHur}^{G,C,(\varphi,\gamma)}_{B,\underline{n}}/B)\to U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}$ is an isomorphism.
@@ 3.3. Brauer group computations
We need to recall some notation from [Loughran2025mallesconjecturebrauergroups] to define this subgroup of the Brauer group. We have provided an alternative description of this subgroup in terms of certain central extensions in §3.3.1 for the reader unfamiliar with [Loughran2025mallesconjecturebrauergroups]. Such a reader may take this alternative description as a definition.
@@
3.3.1. Description of the Brauer group
We will use the non-standard notation $\mathbb{Q}/\mathbb{Z}(1):=\lim_{p\nmid n}\mu_{n}$ to keep the notation light.
@@ 3.3.1. Description of the Brauer group
Consider a central extension of group schemes $1\to\mathbb{Q}/\mathbb{Z}(1)\to E\to G\to 1$ of group schemes over $\mathbb{F}_{q}$. This defines an element $E]\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ by [Loughran2025mallesconjecturebrauergroups, Def. 6.5]. This induces an isomorphism between $BG_{\bar{\mathbb{F}}_{q}}$ and the group of $\mathbb{Q}/\mathbb{Z}(1)$ central extensions up to isomorphism under the Baer sum. This follows by taking the colimit over all $p\nmid n$ of [Loughran2025mallesconjecturebrauergroups, Lem. 6.6,6.7] and the fact that $\operatorname{Br}\mathbb{F}_{q}=0$.
@@
Both statements follows immediately from [Loughran2025mallesconjecturebrauergroups, Lem. 6.9] by taking the colimit over all $p\nmid n$ and writing down the torsors in loc. cit. in terms of cocycles. ∎
3.3.2. Computations
To prove this theorem we will essentially need to do some computations in the derived category $\mathcal{D}(\operatorname{Spec}\mathbb{F}_{q})$ of $\operatorname{Gal}(\bar{\mathbb{F}}_{q}/\mathbb{F}_{q})$-modules. But everything is simple enough that we can give cocycle-theoretic descriptions of all involved objects and will not actually need any facts about derived categories. We leave the argument that the cocycle-theoretic definitions agree with the analogous definitions in the derived category to the interested reader, it follows from a standard bar complex argument.
@@ 3.3.2. Computations
The expression $\psi(\mathrm{Frob}_{q})(\underline{n}+\mathfrak{R}(\tilde{\gamma}))-\alpha_{*}(\eta)(\mathrm{Frob}_{q}^{-1}))\in\mathbb{Q}/\mathbb{Z}$ is equal to
@@ 4.1. Point-counting on Hurwitz spaces
    Consider a chain of subsets $\mathcal{C}_{\mathrm{st}}\subseteq\mathcal{C}\subseteq\mathcal{C}_{G}^{*}$ such that the conjugacy classes in $\mathcal{C}_{\mathrm{st}}$, resp. $\mathcal{C}$, generate $G/Z(G)$, resp. $G$. Let $C\subset G(-1)$ be the subscheme which is the union of elements of $\mathcal{C}$.
@@
    $\mathrm{CHur}^{G,(\varphi,C,\gamma)}_{\underline{n}}(\mathbb{F}_{q})=\#U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}(\mathbb{F}_{q})\#\mathrm{Conf}_{\mathcal{C},\underline{n}}(\mathbb{F}_{q})+O\left(q^{|\underline{n}|-\delta\min_{c\in\mathcal{C}_{\mathrm{st}}(\bar{\mathbb{F}}_{q})}(n_{c})}\right).$ (4.2)
@@
Let us now consider (2). By part (1) and Lemma 3.23 we may assume that the Frobenius fixed components of $\mathrm{CHur}^{G,C,(\varphi,\gamma)}_{\underline{n}}$ are in bijection with $U(G_{\varphi},\mathcal{C})_{\underline{n},\gamma}(\mathbb{F}_{q})$. It thus suffices to show that if $Z\subset\mathrm{CHur}^{G,\varphi}_{\underline{n}}$ is a Frobenius fixed component of a Hurwitz scheme then
@@
Let $\mathcal{C}_{\mathrm{st}}\subset\mathcal{C}\subset\mathcal{C}_{G}^{*},\underline{n}$ be as in Lemma˜4.1(2). Then $\mathrm{CHur}_{\underline{n}}^{G,C,(\varphi,\gamma)}(\mathbb{F}_{q})$ is equal to
$\frac{\#\mathrm{Conf}_{\mathcal{C},\underline{n}}(\mathbb{F}_{q})}{|G^{\mathrm{ab}}(-1)(\mathbb{F}_{q})|}\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}}e(\partial_{\gamma}(\beta)(\varphi))\prod_{c\in\pi_{0}(\mathcal{C})}e(n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta))+O(q^{|\underline{n}|-\delta\min_{c\in\mathcal{C}_{\mathrm{st}}(\bar{\mathbb{F}}_{q})}(n_{c})}).$
@@
$\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}}e\left(\partial_{\gamma}(\beta)(\varphi)+\sum_{c\in\pi_{0}(\mathcal{C})}n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta)))\right)\frac{\#\mathrm{Conf}_{\mathcal{C},\underline{n}}(\mathbb{F}_{q})|\operatorname{H}_{2}(G,\mathcal{C})(\mathbb{F}_{q})|}{|\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\mathbb{F}_{q}}|}.$
@@ 4.2.2. Brauer-Manin pairing
If $P=\infty$ then we have the following explicit formula, where we recall from the abbreviation $\partial_{\gamma}(\beta):=\partial_{c_{\gamma}}(\beta)((\varphi,\gamma))$ from just before Theorem˜3.28.